# Keyed Chaotic Tensor Transformations for Secure and Attributable Neural Inference

### v1

@@ Abstract
This work introduces a novel framework for secure and privacy-preserving neural network inference based on keyed chaotic dynamical transformations. The proposed method applies a deterministic, cryptographically seeded chaotic system to tensors, producing non-invertible, user-specific transformations that enable authenticated inference, tensor-level watermarking, and data attribution. This framework offers a scalable and lightweight alternative to conventional cryptographic techniques, and establishes a new direction for tensor-level security in AI systems.
Note: This is a preliminary preprint released to share the core ideas and invite feedback. Empirical validation and formal security analysis are planned for revisions.
@@ 1 Introduction
This work introduces a framework for applying keyed chaotic dynamical transformations to tensors used in machine learning pipelines, enabling privacy-preserving, traceable, and non-invertible manipulation of data representations. Each transformation is derived from a cryptographically seeded dynamical system, represented as a computation graph that operates deterministically on the input. The transformation can be configured to either preserve task-relevant structure or intentionally obscure it, depending on the application, resulting in representations that are non-invertible, watermarked, and cryptographically bound to a key. These transformations can be applied generally to tensors and are designed to integrate cleanly with existing models, without modifying architecture or requiring retraining.
@@
  • A method for applying deterministic chaotic transformations to tensors using keyed graph dynamical systems
  • A demonstration of how these transformations enable authenticated inference, watermarking, and attribution
  • A discussion of downstream applications including secure data contribution, dynamic ML firewalls, and tensor hash functions
This framework bridges ideas from cryptography, dynamical systems, and representation learning, offering a new lens on how models and data can interact securely. Unlike traditional cryptographic methods, this approach is architecture-agnostic, low-overhead, and enables per-user control over inference contributions. The following sections outline the system design, present empirical results on secure inference and watermark robustness, and explore future directions in secure, chaos-based ML infrastructure.
@@ 2.2 Privacy-Preserving Inference
The approach introduced in this work departs from these paradigms. Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior to or following inference. These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally obfuscated, and task-compatible. While the transformations are reversible by the key holder (e.g., via subtraction or reapplication), they remain opaque to unauthorized parties without knowledge of the underlying system. This enables lightweight, key-authenticated inference, where only correctly transformed inputs yield meaningful results, and offers tensor-level control over privacy, traceability, and secure data contribution—without requiring architectural changes or incurring the computational costs of conventional cryptographic methods.
@@ 3.1 Graph Topology Sampling
Given a cryptographic key $k\in\{0,1\}^{n}$, a pseudorandom seed is derived as
$\text{seed}=\texttt{Hash}(k)\bmod 2^{32},$ (1)
where Hash is a standard cryptographic hash function (e.g., SHA-256), and the result is used to initialize a pseudorandom number generator. This generator is then used to deterministically sample a graph $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
@@ 3.4 Key-Conditioned System Assembly
  • Deterministic: reproducible with the key, without requiring key storage
  • Secure: opaque to observers without the key
  • Flexible: compatible with a range of graph and dynamic structures
@@
3.5 Security Properties
The proposed system provides the following properties:
  • Key-conditioned determinism: Each user’s transformation system is deterministically derived from their cryptographic key, ensuring reproducibility without shared state or explicit key transmission.
  • Input obfuscation: Transformed inputs are non-trivial to interpret without the key, due to both structural scrambling and sensitivity to initialization. Small changes in the key or input yield divergent outputs.
  • Attribution and binding: Only a key holder can produce a given transformation. This allows inference outputs to be linked to specific users or devices, supporting attribution and misuse detection.
  • Opaque trajectories: Without knowledge of the graph topology, weights, and update rule, reverse-engineering the transformation is computationally impractical. The chaotic dynamics ensure exponential divergence from any incorrect reconstruction attempt.
@@ 4.1 Key-Authenticated Neural Network Inference
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a sequence of token embeddings. A user constructs a chaotic transformation system from a cryptographic key $k$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$. The input is scrambled client-side via:
@@
Only if the correct key is used will this inversion yield a valid input. Any incorrectly scrambled input results in faulty or degraded inference, effectively binding model access to possession of the correct key.
@@ 4.3 Data Privacy
where $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ is a deterministic mask generated from a graph dynamical system seeded by the cryptographic key $k$. The scrambled embeddings $\tilde{\mathbf{X}}$ retain the structure and dimensions required for downstream tasks, but are semantically obfuscated and difficult to interpret without knowledge of the system.
@@
This obfuscation protects the user’s input from inspection in transit or at rest, while still enabling standard inference with minimal computational overhead. Furthermore, reverse engineering the obscured tensors is impractical without access to the key generating the chaotic transformation of the tensor hence securing the users raw data.
@@ 4.4 Secure Data Contribution to Shared Models
The proposed transformation framework allows users to submit data obfuscated by a key-conditioned chaotic transformation. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of input embeddings, and let $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ be the mask generated from the user’s key $k$. The transformed data is computed as:
@@
to recover the original data for training. Alternatively, the model could be trained directly on $\tilde{\mathbf{X}}$, treating the transformation as a privacy-preserving encoding layer. However, training directly on obfuscated data may degrade performance depending on the task, and its effectiveness is likely to be application-specific. Further empirical study is needed to determine the conditions under which learning from transformed data remains effective.
@@ 5 Discussion
While the system provides functional security properties such as key-conditioned masking, user attribution, and trajectory-based unpredictability, it does not yet offer formal cryptographic guarantees. In particular, the transformation is reversible by design for authorized users, and may be vulnerable to chosen-input or model inversion attacks in some threat models. As such, it is best viewed as a complementary privacy mechanism—sitting between full encryption and raw data sharing—and may be combined with other techniques (e.g., differential privacy, TEEs, encrypted transport) in high-risk applications.
@@ 6 Conclusion
This work introduces a novel framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning. By leveraging deterministic chaos seeded from cryptographic keys, the method enables input obfuscation, output watermarking, and authenticated inference without modifying the underlying model. The approach is lightweight, task-compatible, and suitable for a range of real-world applications where privacy, attribution, or secure contribution is essential. As the need for secure AI deployment grows, this method offers a new building block at the interface of dynamical systems, cryptography, and neural computation.

### v2

# Keyed Chaotic Masking: A Functional Privacy Framework for Neural Inference

@@ Abstract
This work introduces a lightweight framework for privacy-preserving neural network inference based on keyed chaotic masking a deterministic, user-specific obfuscation method derived from cryptographically seeded chaotic dynamical systems. The approach applies masks to input and output tensors using key-conditioned graph dynamics, enabling authenticated inference, user attribution, and soft output watermarking without modifying model architectures. While the underlying chaotic system used to generate each mask is not analytically invertible, the masking operation itself is algebraically reversible by authorized key holders, offering functional privacy without formal cryptographic guarantees. Unlike traditional encryption or secure multi-party computation, this method operates in continuous space and imposes minimal computational overhead. We describe the construction of the masking system, including graph sampling, dynamical rule selection, and chaos diagnostics. Applications include privacy-preserving inference, secure data contribution, and per-user watermarking in shared model pipelines. This framework offers a practical and modular building block for user-controlled privacy in modern AI systems.
Note: This is a preliminary preprint released to share the core ideas and invite early feedback. The proposed framework uses cryptographically derived keys (e.g., via ECDH and HKDF) to seed user-specific transformations, and inherits security from these primitives. However, the chaotic transformation itself is not a formally proven cryptographic primitive. It should be viewed as a functional privacy mechanism, offering deterministic obfuscation and user attribution. Future work will explore its robustness under adversarial threat models and its relationship to formal notions of security.
@@ 1 Introduction
This work introduces a framework for applying keyed chaotic dynamical transformations to tensors used in machine learning pipelines, enabling privacy-preserving, traceable, and non-invertible manipulation of data representations. Each transformation is derived from a cryptographically seeded dynamical system, represented as a computation graph that operates deterministically on the input. The transformation can be configured to either preserve task-relevant structure or intentionally obscure it, depending on the application, resulting in representations that are effectively non-invertible due to chaotic divergence, watermarked, and cryptographically bound to a key. These transformations can be applied generally to tensors and are designed to integrate cleanly with existing models, without modifying architecture or requiring retraining.
@@
  • A framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning pipelines, enabling user-specific, deterministic, and lightweight obfuscation of data.
  • A construction method for cryptographically seeded graph-based dynamical systems, including deterministic graph sampling, node-level chaotic update rules, and Lyapunov-based validation.
  • A conceptual demonstration of the method’s utility in input obfuscation, output watermarking, and authenticated inference, showing how these transformations support data privacy, user attribution, and secure access.
  • An exploration of downstream use cases, including privacy-preserving inference, federated data contribution, and soft watermarking for multi-user systems—highlighting the compatibility of the approach with existing model architectures.
This framework bridges ideas from cryptography, dynamical systems, and representation learning, offering a new lens on how models and data can interact securely. Unlike traditional cryptographic methods, this approach is architecture-agnostic, low-overhead, and enables per-user control over inference contributions. The following sections outline the system design, propose future experiments on secure inference and watermark robustness, and explore directions for chaos-based ML infrastructure.
@@ 2.2 Privacy-Preserving Inference
The approach introduced in this work departs from these paradigms. Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior to or following inference. These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally transformed, and task-compatible (i.e., preserving shape and structure for downstream model compatibility). While the masking operation is algebraically reversible by the key holder (e.g. via subtraction), the underlying chaotic system used to generate the mask is not analytically invertible. As a result, the overall transformation remains opaque to unauthorized parties. The chaotic nature of the transformation ensures exponential sensitivity to small changes in initial state or key, rendering inversion impractical without full system knowledge—even though the transformation is technically deterministic. This enables lightweight, key-authenticated inference, where only correctly transformed inputs yield meaningful results, and offers tensor-level control over privacy, traceability, and secure data contribution—without requiring architectural changes or incurring the computational costs of conventional cryptographic methods.
@@ 3.1 Graph Topology Sampling
Given a cryptographically derived key $k$, a pseudorandom byte stream is generated using a deterministic key derivation function (e.g., HKDF with SHA-256). This stream is used to seed a pseudorandom number generator for deterministic graph sampling and system assembly. When required for compatibility with legacy APIs, a 32-bit integer seed may be extracted from the first four bytes of the stream.
The resulting generator is used to deterministically sample a graph $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
@@ 3.4 Key-Conditioned System Assembly
  • Deterministic: reproducible with the key
  • Key-obscured: opaque to observers without the key
  • Composable: the system can be integrated into existing pipelines without modifying model architecture or training objectives
@@ 4 Applications
The applications described in this section are conceptual demonstrations based on the properties of the transformation system. While they are technically implementable and initial implementation confirms Lyapunov divergence across key configurations; empirical evaluations of inference compatibility, robustness to key error, and resistance to inversion attacks are ongoing. In all applications, we assume a shared key is either held by the user or securely negotiated with the server via key exchange (e.g., ECDH). The transformation can be applied on either side of the interaction, depending on the privacy model and deployment scenario.
@@ 4.1 Key-Authenticated Neural Network Inference
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a sequence of token embeddings. A user constructs a chaotic transformation system from a cryptographic key $k$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$. The input is masked client-side via:
@@
Only if the correct key is used will this inversion yield a valid input. Any incorrectly masked input results in faulty or degraded inference, effectively binding model access to possession of the correct key.
@@ 4.3 Data Privacy
where $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ is a deterministic mask generated from a graph dynamical system seeded by the cryptographic key $k$. The masked embeddings $\tilde{\mathbf{X}}$ retain the structure and dimensions required for downstream tasks, but are semantically transformed and difficult to interpret without knowledge of the system.
@@
This obfuscation protects the user’s input from inspection in transit or at rest, while still enabling standard inference with minimal computational overhead. Furthermore, reverse engineering the obscured tensors is impractical without access to the full chaotic transformation, as the dynamics induce exponential divergence with respect to both key and input, making inversion intractable without the key.
@@ 4.4 Secure Data Contribution to Shared Models
The proposed transformation framework allows users to submit data masked by a key-conditioned chaotic transformation. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of input embeddings, and let $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ be the mask generated from the user’s key $k$. The transformed data is computed as:
@@
to recover the original data for training. Alternatively, the model could be trained directly on $\tilde{\mathbf{X}}$, treating the transformation as a privacy-preserving encoding layer. However, training directly on transformed data may degrade performance depending on the task, and its effectiveness is likely to be application-specific. Further empirical study is needed to determine the conditions under which learning from transformed data remains effective.
@@ 5 Discussion
While the system is designed to offer functional security properties such as key-conditioned masking and trajectory-based unpredictability, these remain to be empirically evaluated in practical deployments. In particular, the transformation is reversible by design for authorized users, and may be vulnerable to chosen-input or model inversion attacks in some threat models. As such, it is best viewed as a complementary privacy mechanism—sitting between full encryption and raw data sharing—and may be combined with other techniques (e.g., differential privacy, TEEs, encrypted transport) in high-risk applications.
@@ 6 Conclusion
This work introduces a novel framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning. By leveraging deterministic chaos seeded from cryptographic keys, the method enables input obfuscation, output watermarking, and authenticated inference without modifying the underlying model. The approach is lightweight, task-compatible, and suitable for a range of real-world applications where privacy, attribution, or secure contribution is essential. As the need for secure AI deployment grows, this method offers a new building block at the interface of dynamical systems, cryptography, and neural computation. Please note that this preprint version is released to share the concept and invite community feedback. Validation experiments and security evaluations are ongoing and will appear in subsequent revisions.